OpenStudy (fanduekisses):

How to prove cos4x+cos2x=2-2sin^2 2x-2sin^2x ? I'm confused :'(

OpenStudy (whpalmer4):

\[\cos 4x + \cos 2x = 2 - 2\sin^2 2x - 2\sin^2 x\]Get out your trig identity crib sheet. Do you see any identities you can use to decompose \(\cos 4x,~\cos 2x,~\sin 2x\) into various combinations of \(\sin x\) and \(\cos x\)?

OpenStudy (fanduekisses):

cos 2x, the double angle one?

OpenStudy (fanduekisses):

idk about the cos 4x :(

OpenStudy (whpalmer4):

that would be one

OpenStudy (whpalmer4):

can't you use the double angle one twice?

OpenStudy (fanduekisses):

oh ok ill try it

OpenStudy (whpalmer4):

you can do it with a substitution like this: \[u = 2x\] \[\cos 4x = \cos 2u\]\[\cos 2u = ...\] then undo the substitution

OpenStudy (fanduekisses):

so cos(2x)= cos4x ?

OpenStudy (whpalmer4):

No... I'm going to do a different identity as a demonstration. \[\sin 2x = 2 \sin x \cos x\]right?

OpenStudy (fanduekisses):

\[\cos2(2x)+1-2\sin ^{2}\]

OpenStudy (fanduekisses):

ok

OpenStudy (whpalmer4):

Say we had \[\sin 4x\]and wanted to decompose it into \(\sin x\) and \(\cos x\) we could say \(u = 2x\) and rewrite our expression as \[\sin 4x = \sin 2u\] any question so far?

OpenStudy (fanduekisses):

ok

OpenStudy (whpalmer4):

Then, we use the identity: \[\sin 2u = 2\sin u \cos u\]Now we undo the substitution, giving us \[\sin 2u =\sin 4x = 2\sin 2x \cos 2x\] But \(\sin 2x = 2 \sin x \cos x\) so we can rewrite that as \[\sin 4x = 2(2\sin x \cos x)\cos 2x\]and then we can use one of the identities for \(\cos 2x\) to split it further

OpenStudy (whpalmer4):

depending on what we were trying to do, we would make our selection from the 3 different choices.

OpenStudy (whpalmer4):

another way to handle the \(\cos 4x\) might be to use the cosine of a sum of angles identity: \[\cos(u+v) = \cos u \cos v - \sin u \sin v\]with \(u = 2x\) and \(v = 2x\)

OpenStudy (fanduekisses):

oh ok I like the sum of angles identity better :)

OpenStudy (fanduekisses):

so cos(2x+2x)= cos2xcosx2x-sin2xsinx2x

OpenStudy (fanduekisses):

oops disregard the extra x's

OpenStudy (whpalmer4):

well, you have some extra x's in there... \[\cos 4x = \cos (2x+2x) = \cos 2x \cos 2x - \sin 2x \sin 2x\] And then you'd want to expand \(\cos 2x\) and \(\sin 2x\)... Might be worth investigating what you're going to get from the other terms before doing so, just to give you an idea of what you might want to choose for the \(\cos 2x\) expansion

OpenStudy (fanduekisses):

so =\[2cox ^{2}x-1(2cox^{2}-1)-(2sinxcosx)\]

OpenStudy (whpalmer4):

if you put a \ before cos or sin, you'll get the proper formatting...

OpenStudy (fanduekisses):

" / " like a fractions?

OpenStudy (whpalmer4):

also, assuming I interpreted what you wrote as you intended it, it isn't correct... no, backslash, look above the "return" key on a US-style keyboard

OpenStudy (fanduekisses):

what does it do? \

OpenStudy (whpalmer4):

the difference is \[cos x\]vs.\[\cos x\] I typed "\ [ c o s x \ ]" for the first one and "\ [ \ c o s x \ ]" for the second one (without putting an extra space between each character, of course)

OpenStudy (whpalmer4):

If you see some interesting formatting and want to know how it is done, you can click and drag across the formula, then right-click and select Show Math As -> Tex Commands" and you'll get a little popup window that shows the code.

OpenStudy (whpalmer4):

but that's not really important to your problem...

OpenStudy (fanduekisses):

ok

OpenStudy (fanduekisses):

so was my answer correct?

OpenStudy (fanduekisses):

what do I do next

OpenStudy (whpalmer4):

I'm not sure what "your answer" was...

OpenStudy (whpalmer4):

but why don't you try expanding the stuff on the right side first...it's a bit simpler, and it will give us some clues as to how we should choose to expand the left side to make it match

OpenStudy (fanduekisses):

ok, oh don't know where to start

OpenStudy (fanduekisses):

ok so 2-2sin^2 2x-2sin^2x

OpenStudy (whpalmer4):

\[\cos 4x + \cos 2x = 2 - 2\sin^2 2x - 2\sin^2 x\] Let's expand the right hand side. \[2 - 2\sin^2 2x - 2\sin^2 x\] can you rewrite that with the expansion of \(\sin 2x\) in place?

OpenStudy (fanduekisses):

sorry, I am back

OpenStudy (whpalmer4):

No problem...

OpenStudy (fanduekisses):

like 2sinxcosx

OpenStudy (whpalmer4):

could you at least put spaces in where they belong, please? 2 sin x cos x

OpenStudy (fanduekisses):

ok :)

OpenStudy (whpalmer4):

or use parentheses around the arguments, like this: \[\cos (2 x)+\cos (4 x)=2 -2 \sin ^2(x)-2 (2 \sin (x) \cos (x))^2\]\[\cos (2 x)+\cos (4 x)=2-2 \sin ^2(x)-4 \sin^2 (x) \cos^2 (x)\]

OpenStudy (whpalmer4):

look at the left two terms on the right side of the equation. Do you see anything you might do there?

OpenStudy (fanduekisses):

\[2(1-\sin^{2})\]

OpenStudy (fanduekisses):

factor?

OpenStudy (whpalmer4):

you need to drag your arguments along for the ride...

OpenStudy (fanduekisses):

:'( I'm just so confused right now

OpenStudy (whpalmer4):

sin of what?!? is that \[2(1-\sin^2x)\]or\[2(1-\sin^2 y)\]or\[2(1-\sin^2(3x^2+4x+5))\]or anything else I might imagine?

OpenStudy (whpalmer4):

no, it's \[2(1-\sin^2x)\]Details matter, especially when doing these annoying proofs that are full of opportunities for error! You must be fastidious in your work... I'm sorry, but I have to leave you now to deal with my other obligations today. Perhaps my friend @mathmale can help you out when he's available.